# Sample Selection Bias in Machine Learning for Healthcare

### v1

@@ 5.2 Analysis with the Synthetic Data
Here, we present a comparative study of various techniques for handling SSB with the synthetic dataset. Since the performance could vary across different event rates, selection rates or dataset sizes, first we present a summary of the performance across different settings using boxplots in Fig. 4, and subsequent subsections study the effect of the selection/event rates and dataset sizes. We consider the event rates and the non-selection rates from [5, 10, 20, 30, 40]%, and dataset sizes from [1000, 2000, 3000, 4000, 5000] [^6] For a dataset size of 1000 with a non-selection rate and an event rate of 10% each, it implies there are 100 non-selected patients and 90 of the remaining 900 have positive outcomes while the rest has negative outcomes., where each setting is repeated and averaged over 10 seeds. We observe large variations in the performance of all the techniques due to diverse settings, where extreme cases like low event rates, low non-selection rates or small dataset sizes could lead to a large drop in the performance.
@@ 5.2.1 Effect of the Non-selection Rates and the Event Rates
To study the effect of the non-selection rates and the event rates on the performance of different techniques to address SSB, we consider the event rates and the non-selection rates from [5, 10, 20, 30, 40]%, and present a summary of the performance averaged over dataset sizes from [1000, 2000, 3000, 4000, 5000]. For the sake of clarity, we consider the proposed techniques and two best baselines, i.e., DANN and KMM, along with the naive techniques, and present differences in their performance from Oracle as heatmaps in Fig. 5.
@@
From Fig. 5, it is clear that overall MT-Net and T-Net perform better than the baselines as the difference with Oracle is lesser than the baselines. The relative performance of Naive technique drops up to 22% for a low non-selection rate and a high event rate, while the lowest drop occurs at a high non-selection rate and a low event rate. This underlies the importance of addressing SSB in machine learning for healthcare as machine learning algorithms showing high performance may not show similar performance when deployed due to SSB. It is interesting to note that for some settings, like a high non-selection rate and a low event rate, all four techniques perform closely to Oracle and then the relative performance decreases as the non-selection rate decreases or the event rate increases, where the baselines show more drop in the performance as compared with the proposed techniques. The worst performance occurs at a low non-selection rate and a high event rate for all techniques, except DANN which shows a slightly different pattern at an event rate of 0.3 and 04. Since we keep the total number of data points fixed while we change one of the two rates, the pattern of a decrease in the relative performance with a decrease in the non-selection rate can be accounted for less available data to address SSB as well as an increase in imbalance for non-selection rate. On the other hand, the pattern of a decrease in the relative performance with an increase in the event rate, i.e., a decrease in class imbalance, for the event rate across all the techniques is a bit counter-intuitive, as generally, the performance improves with improvement in class imbalance. Further analysing, we observe that the performance of individual techniques does increase with an increase in the event rate, however, the relative performance presented in the heatmaps shows a decrease because Oracle shows less variability with event rate than the rest of the techniques resulting in a decrease in the performance with an increase in event rate. These results also show a consistency of the proposed techniques across different settings as compared with the baselines. Overall, Naive performs the worst and MT-Net performs the best to address SSB.
@@ 5.3 Analysis with the Semi-synthetic Data
On analysing the performance of Naive technique as compared with Oracle, we observe an average drop of two percent and three percent for COVID-19 and Diabetes, respectively. Oracle performs the best in terms of median performance for COVID-19, however, for Diabetes dataset, KLIEP and DANN show performance similar to Oracle. Oracle is closely followed by T-Net and MT-Net, however, MT-Net shows the least variation in the performance among all the techniques across all the settings. The smaller variation in the performance of MT-Net occurs because it has access to included as well as excluded data points, resulting in shared learning between the tasks, which is helpful with limited data settings. Similar to the synthetic dataset, Imputation shows the worst performance both in terms of mean performance as well as variation in the performance. Interestingly, the performance of KMM and Imputation for COVID-19 and MT-Naive, DANN and Imputation for Diabetes falls below Naive method which indicates that the bias handling negatively impacts the performance. We study the effect of the non-selection rates, event rates and dataset sizes in the following subsections.
@@ 5.3.1 Effect of the Non-selection Rates and the Event Rates
To study the effect of the non-selection rates and the event rates on the performance of different techniques to address SSB, we consider the event rates and the non-selection rates from [5, 10, 20, 30, 40]%, and present a summary of the performance averaged over dataset sizes from [1000, 2000, 5000, 10,000, 15,000/25,000] for COVID-19/Diabetes datasets. For the sake of clarity, we consider the proposed techniques and two best baselines, i.e., DANN and KLIEP for COVID-19 and KMM and KLIEP for Diabetes datasets, along with the naive techniques, and present differences in their performance from Oracle as heatmaps in Fig. 8.
@@
From Fig. 8, it is clear that overall MT-Net and T-Net perform better than the baselines as the difference with Oracle is lesser than the baselines. The relative performance of Naive technique drops more than 6% for COVID-19 and more than 10% for Diabetes datasets. As noted earlier, this underlies the importance of addressing SSB in machine learning for healthcare. Similar to the synthetic dataset, overall, the performance difference from Oracle decreases as the event rate increases or the non-selection rate decreases, however, the non-selection rate has more effect on the performance than the event rate. Interestingly, all the techniques show improvements over Oracle in certain settings, as indicated by negative values in heatmaps, and in those settings, MT-Naive performs the best. This happens when the non-selected patients lie on the correct side of the decision boundary and away from the decision boundary that they do not have any role in deciding the decision boundary. In such cases, the naive techniques will make correct predictions even if they have not seen those patients during the training, while all the bias handling techniques would lead to some loss of performance. Once again, these results show a consistency of the proposed techniques across different settings as compared with the baselines.
@@ 5.3.2 Effect of Dataset Size
From the figure, we observe that MT-Net and T-Net techniques perform close to Oracle, which is the desired behaviour when we are addressing SSB in the dataset. Similar to the synthetic dataset, Imputation performs the worst for both the datasets, and Oracle performs the best on Diabetes dataset but for COVID-19 dataset MT-Naive performs the best. For COVID-19 dataset, KMM and Imputation perform worse than Naive, while for Diabetes dataset, MT-Naive, Imputation and DANN perform worse than Naive. This indicates the bias correction techniques fail and negatively impact the performance. As observed earlier, MT-Net is more effective than T-Net for smaller datasets due to access to included and excluded patients and from sharing of information. The proposed techniques perform consistently across settings and datasets, while the baselines show huge fluctuations, for example, MT-Naive outperforms the rest for COVID-19 but falls even below Naive for Diabetes. Similarly, KLIEP shows good performance for smaller datasets as observed for COVID-19 and Diabetes but lagged for the synthetic dataset. Overall, as expected, the performance of all the techniques for addressing SSB improves with an increase in the dataset size.
5.4 Effect of SSB on the Non-selected Population
We analyse the performance of the best SSB handling baselines to study their performance for the selected, the non-selected and all the patients of the study with the synthetic and semi-synthetic datasets, and present results in Fig. 10.
@@ 5.4 Effect of SSB on the Non-selected Population
5.5 Analysis of the TPI Approach
Here, we analyse the performance of the proposed MT-Net and T-Net techniques, measured using AUC, to identify the target population with the synthetic, COVID-19 and Diabetes datasets, and present results in Fig. 11.
@@ 6 Discussion
The existing techniques, borrowed from other machine learning areas like domain-adaption, are mostly based on the idea of bias correction and try to match distributions between the study population and the target population, leading to a loss of predictive performance of the machine learning algorithms. To address this issue, we propose a new research direction to address SSB based on the TPI, rather than bias correction. Here, an algorithm would focus first on identifying the target subpopulation representative of the study population and seen by the algorithm during the training time, and then making predictions for the identified subpopulation, while deferring the rest of the patients to clinicians. This approach bears resemblance to uncertainty quantification where predictions are generated only for those patients where an algorithm is confident otherwise the patients are referred to a clinician [38, 12]. Uncertainty quantification and TPI can also be regarded as subcategories within the broader framework of machine learning with rejection, where algorithms refrain from making predictions when they are likely to err [29]. The proposed approach is motivated by healthcare where certain patient populations fail to complete a study and the target population would contain any patients, including patient groups who failed to complete the study. This setting is unlike domain adaptation where the target dataset is supposed to be different from the source and the algorithms are supposed to work with the target dataset. Additionally, compared to the naive algorithms, the proposed approach can utilise the non-selected patients for training machine learning algorithms which could be helpful for limited data settings.
@@
The proposed techniques make predictions only for the target subpopulation representative of the study population, i.e., the selected patients. However, we want to argue that the non-random sample selection is inevitable in healthcare, for example, all the clinical studies have strict inclusion and exclusion criteria (say adult patients) implying that the results apply to the restricted population (only to adult patients). So, as long as the study and the target population are similar, the bias does not occur and the results/algorithms are still useful. It is better to make correct predictions for subpopulations than to make biased predictions for certain subpopulations when considering the entire population. Thus, the proposed SSB handling approach ensures robust and reliable predictions within the defined subpopulation, contributing to more accurate and effective healthcare decision-making.

### v2

@@ 5.2 Analysis with the Synthetic Data
Here, we present a comparative study of various techniques for handling SSB with the synthetic dataset. Since the performance could vary across different event rates, selection rates or dataset sizes, first we present a summary of the performance across different settings using boxplots in Fig. 4, and subsequent subsections study the effect of the selection/event rates and dataset sizes. We consider the event rates and the non-selection rates from [5, 10, 20, 30, 40]as we don’t expect large non-selection rates or event rates in real-world datasets, and dataset sizes from [1000, 2000, 3000, 4000, 5000] [^6] For a dataset size of 1000 with a non-selection rate and an event rate of 10% each, it implies there are 100 non-selected patients and 900 selected patients. 90 Of the 900 selected patients have positive outcomes while the rest 810 has negative outcomes., where each setting is repeated and averaged over 10 seeds. We observe large variations in the performance of all the techniques due to diverse settings, where extreme cases like low event rates, low non-selection rates or small dataset sizes could lead to a large drop in the performance.
@@ 5.2.1 Effect of the Non-selection Rates and the Event Rates
To study the effect of the non-selection rates and the event rates on the performance of different techniques to address SSB, we consider the event rates and the non-selection rates from [5, 10, 20, 30, 40]%, and present a summary of the performance averaged over dataset sizes from [1000, 2000, 3000, 4000, 5000]. For the sake of clarity, we consider the proposed techniques and two best baselines, i.e., DANN and KMM, along with the naive techniques, and present absolute differences in their performance from Oracle as heatmaps in Fig. 5 because Oracle presents an ideal case and unbiased model.
@@
From Fig. 5, it is clear that overall MT-Net and T-Net perform better than the baselines as the difference with Oracle is lesser than the baselines. The relative performance of Naive technique drops up to 22% for a low non-selection rate and a high event rate, while the lowest drop occurs at a high non-selection rate and a low event rate. This underlies the importance of addressing SSB in machine learning for healthcare as machine learning algorithms showing high performance may not show similar performance when deployed due to SSB. It is interesting to note that for some settings, like a high non-selection rate and a low event rate, all four techniques perform closely to Oracle and then the relative performance decreases as the non-selection rate decreases or the event rate increases. The baselines show more drop in the performance as compared with the proposed techniques. The worst performance occurs at a low non-selection rate and a high event rate for all techniques, except DANN which shows a slightly different pattern at an event rate of 0.3 and 04. Since we keep the total number of data points fixed while we change one of the two rates, the pattern of a decrease in the relative performance with a decrease in the non-selection rate can be accounted for less available data to address SSB as well as an increase in imbalance for non-selection rate. On the other hand, the pattern of a decrease in the relative performance with an increase in the event rate, i.e., a decrease in class imbalance, for the event rate across all the techniques is a bit counter-intuitive, as generally, the performance improves with improvement in class imbalance. Further analysing, we observe that the performance of individual techniques does increase with an increase in the event rate, however, the relative performance presented in the heatmaps shows a decrease because Oracle shows less variability with event rate than the rest of the techniques resulting in a decrease in the performance with an increase in event rate. These results also show a consistency of the proposed techniques across different settings as compared with the baselines. Overall, Naive performs the worst and MT-Net performs the best to address SSB.
@@ 5.3 Analysis with the Semi-synthetic Data
On analysing the performance of Naive technique as compared with Oracle, we observe an average drop of two percent and three percent for COVID-19 and Diabetes, respectively. Moreover, while Oracle performs the best in terms of median performance for the COVID-19 and Diabetes datasets, we observe some baselines performing close to Oracle. For example, in the Diabetes dataset, KLIEP and DANN show performance like Oracle. The small drop in average performance across different settings and the good performance of some baselines can be attributed to combination of two factors. First, except for Oracle and the proposed techniques, the remaining techniques make predictions for all patients (i.e., included and excluded). Second, for certain combinations of non-selection rates and event rates, the non-selected patients do not influence the decision function of the prediction model (This happens when the non-selected patients lie on the correct side of the decision boundary and away from the decision boundary that they do not have any role in deciding the decision boundary.). Consequently, all techniques that make predictions for all patients consistently make correct predictions for the non-selected patients, which improves their overall performance for all patients. This is also reflected in Fig. 7, where the maximum performance of some baselines is even higher than that of Oracle. MT-Net shows the least variation in the performance among all the techniques across all the settings. The smaller variation in the performance of MT-Net occurs because it has access to included as well as excluded data points, resulting in shared learning between the tasks, which is helpful with limited data settings. Similar to the synthetic dataset, Imputation shows the worst performance both in terms of mean performance as well as variation in the performance. Interestingly, the performance of KMM and Imputation for COVID-19, and MT-Naive, DANN and Imputation for Diabetes falls below Naive method which indicates that the bias handling negatively impacts the performance. We study the effect of the non-selection rates, event rates and dataset sizes in the following subsections.
@@ 5.3.1 Effect of the Non-selection Rates and the Event Rates
To study the effect of the non-selection rates and the event rates on the performance of different techniques to address SSB, we consider the event rates and the non-selection rates from [5, 10, 20, 30, 40]%, and present a summary of the performance averaged over dataset sizes from [1000, 2000, 5000, 10,000, 15,000/25,000] for COVID-19/Diabetes datasets. For the sake of clarity, we consider the proposed techniques and two best baselines, i.e., DANN and KLIEP for COVID-19 and KMM and KLIEP for Diabetes datasets, along with the naive techniques. We present results as absolute differences in their performance from Oracle as heatmaps in Fig. 8 because Oracle presents an ideal case and unbiased model.
@@
From Fig. 8, it is clear that overall MT-Net and T-Net perform better than the baselines as the difference with Oracle is lesser than the baselines. The relative performance of Naive technique drops more than 6% for COVID-19 and more than 10% for Diabetes datasets. As noted earlier, this underlies the importance of addressing SSB in machine learning for healthcare. Similar to the synthetic dataset, overall, the performance difference from Oracle decreases as the event rate increases or the non-selection rate decreases, however, the non-selection rate has more effect on the performance than the event rate.
We observe that MT-Naive and DANN exhibit the highest deviations from Oracle for the COVID-19, with values of 0.227 and 0.166, respectively. For the Diabetes, MT-Naive and KLIEP show the highest deviations, at 0.135 and 0.080, respectively. In contrast, T-Net performs the best across all settings, with maximum deviations of only 0.002 for COVID-19 and 0.005 for Diabetes. Similarly, MT-Net also shows lower deviations as compared to baselines. Once again, these results show a consistency of the proposed techniques across different settings as compared with the baselines.
@@ 5.3.2 Effect of Dataset Size
From the figure, we observe that MT-Net and T-Net techniques perform close to Oracle, which is the desired behaviour when we are addressing SSB in the dataset. Similar to the synthetic dataset, Imputation performs the worst for both the datasets, and Oracle performs the best on Diabetes dataset but for COVID-19 dataset MT-Naive performs the best. For COVID-19 dataset, KMM and Imputation perform worse than Naive, while for Diabetes dataset, MT-Naive, Imputation and DANN perform worse than Naive. This indicates the bias correction techniques fail and negatively impact the performance.
As observed earlier, MT-Net is more effective than T-Net for smaller datasets due to access to included and excluded patients and from sharing of information. The proposed techniques perform consistently across settings and datasets, while the baselines show huge fluctuations, for example, MT-Naive outperforms the rest for COVID-19 but falls even below Naive for Diabetes. Similarly, KLIEP shows good performance for smaller datasets as observed for COVID-19 and Diabetes but lagged for the synthetic dataset. Overall, as expected, the performance of all the techniques for addressing SSB improves with an increase in the dataset size.
5.4 Effect of SSB on the Non-selected Subpopulation
We analyse the performance of the best SSB handling baselines to study their performance for the selected, the non-selected and all the patients of the study with the synthetic and semi-synthetic datasets, and present results in Fig. 10777We can’t perform such an analysis for the proposed techniques as they make predictions only for the selected subpopulation..
@@ 5.4 Effect of SSB on the Non-selected Subpopulation
5.5 Ablation of TPI Approach
Here, we analyse MT-Net and T-Net’s ability to identify the target population with the synthetic, COVID-19 and Diabetes datasets, measured using AUC. The results are presented in Fig. 11.
@@ 6 Discussion
The existing techniques, borrowed from other machine learning areas like domain-adaption, are mostly based on the idea of bias correction and try to match distributions between the study population and the target population, leading to a loss of predictive performance of the machine learning algorithms. To address this issue, we propose a new research direction to address SSB based on the TPI, rather than bias correction. Here, an algorithm would focus first on identifying the target subpopulation representative of the study population and seen by the algorithm during the training time, and then making predictions for the identified subpopulation, while deferring the rest of the patients to clinicians. Thus, the proposed approach avoids the loss of predictive power due to alignment of the study and the target population distributions as is reflected by consistent superior performance across settings, including varying non-selection rates, event rates, dataset sizes as well as across multiple datasets. This approach bears resemblance to uncertainty quantification where predictions are generated only for those patients where an algorithm is confident otherwise the patients are referred to a clinician [41, 12] which is essential for the safe application of machine learning [28, 47]. Uncertainty quantification and TPI can also be regarded as subcategories within the broader framework of machine learning with rejection, where algorithms refrain from making predictions when they are likely to err [32]. The proposed approach is motivated by healthcare where certain patient populations fail to complete a study and the target population would contain any patients, including patient groups who failed to complete the study. This setting is unlike domain adaptation where the target dataset is supposed to be different from the source and the algorithms are supposed to work with the target dataset. Additionally, compared to the naive algorithms, the proposed approach can utilise the non-selected patients for training machine learning algorithms which could be helpful for limited data settings.
@@
The proposed techniques make predictions only for the target subpopulation representative of the study population, i.e., the selected patients. However, we argue that non-random sample selection is inevitable in healthcare. For example, clinical studies often have strict inclusion and exclusion criteria (e.g., only adult patients), meaning the results apply to the restricted population (only adult patients). Therefore, as long as the study and target population align, bias does not occur, and the algorithms remain useful. It is better to make accurate predictions for subpopulations than to make biased predictions for certain subpopulations when considering the entire population. Thus, our SSB handling approach ensures robust and reliable predictions within the defined subpopulation, contributing to more accurate and effective healthcare decision-making.
The proposed techniques for selective predictions may raise ethical concerns if the non-represented patients in the study population belong to marginalised subgroups, thus exacerbating existing social and health inequalities. There are three potential approaches to addressing this issue, as discussed in [63, 25]: (i) not using algorithms at all – which also presents an ethical dilemma, as not using algorithms could harm subpopulations that might benefit from their application, (ii) employing bias-correction algorithms, which risk making harmful decisions for non-represented subpopulations, and (iii) using selective predictions and referring non-selected patients to clinicians. While both options one and two present ethical concerns, option three – selective predictions – is a potentially viable approach emerging in fields like bioethics [63]. By referring non-represented subpopulations to clinicians, their care is not negatively affected, while selective predictions could benefit certain subpopulations. While our focus is on developing methodology, we believe that the use of selective predictions should be context-dependent and involve input from the relevant stakeholders.